### v1

@@ 1 Introduction
Computer-assisted diagnosis (CAD) helps specialists such as radiologists and surgeons to interpret medical images, including magnetic resonance imaging (MRI), computed tomography (CT), and X-ray images. The application of deep learning techniques to medical images [1, 2, 3, 4] has yielded increasingly satisfactory results, making it a popular research focus, especially in fracture detection [5, 6, 7].
You Only Look Once (YOLO) series [8, 9, 10, 11, 12, 13, 14, 15, 16] are the main neural networks for real-time object detection task, widely employed in fracture detection [17, 18, 19]. Wrist fractures in children are more common cases and the GRAZPEDWRI-DX dataset [20] provides 20,327 X-ray images of pediatric wrist trauma that can be used in fracture detection tasks. Research [21] first used the YOLOv8 [16] model for fracture detection on this dataset. Since attention mechanisms [22, 23, 24, 25] have excellent results in enhancing the performance of neural network models, Chien et al. achieved the state-of-the-art (SOTA) performance by incorporating different attention mechanisms into the YOLOv8 model.
With the presentation of YOLOv9 [26], which achieved remarkable model performance on the MS COCO 2017 [27] benchmark dataset, this paper first trained the YOLOv9 model on the GRAZPEDWRI-DX dataset and obtained the SOTA performance, as shown in Fig 1.
@@
Figure 1: Comparisons of fracture detection models on the GRAZPEDWRI-DX dataset. In terms of accuracy, our models outperform all previous models at the state-of-the-art level.
@@ 2 Related Works
In the field of object detection task, detectors typically employ either one-stage or two-stage algorithms. Compared to two-stage object detectors, the models of YOLO series offer a more balanced combination of accuracy and inference speed, making them suitable for deployment on mobile computing platforms for medical image recognition. Son et al. [28] utilized YOLOv4 [9] and U-Net [29] as auxiliary diagnostic tools to assist dentists in identifying mandibular fractures without resorting to cone beam computed tomography (CBCT). Jeon et al. [30] employed YOLOv4 [9] to aid surgeons in diagnosing trauma by detecting the fracture and mapping it onto a 3D reconstructed bone image, providing a clear display of the fracture region through the red mask overlaid on the 3D bone image. Hržić et al. [18] employed the YOLOv4 [9] model for fracture detection on the GRAZPEDWRI-DX dataset [20], which was the first to demonstrate that the models of YOLO series can assist radiologists in more accurately predicting wrist injuries in children on X-ray images. Ahmed et al. [31] demonstrated the potential of one-stage algorithm models in enhancing the accuracy of diagnosis on pediatric wrist X-ray images by employing YOLOv5 [12], YOLOv6 [13], YOLOv7 [15], and YOLOv8 [16] models for wrist anomaly detection, respectively. Warin et al. [32] utilized the YOLOv5 [12] model to detect mandibular fractures in panoramic X-ray images, demonstrating the ability of the YOLOv5 model to recognize mandibular fractures at an expert level. Gaikwad et al. [33] applied the YOLOv5 [12] model to detect major and minor fractures of the C1 to C7 vertebrae, achieving an accuracy rate of 89%. Zou et al. [34] investigated various fracture morphologies throughout the body, including angle fractures, normal fractures, line fractures, and disoriented angle fractures. They integrated the YOLOv7 [15] model with the attention mechanism [35], achieving the superior performance on the FracAtlas [36] dataset. Samothai et al. [17] demonstrated that the YOLOX [10] model exhibits faster convergence speed and higher accuracy than YOLOR [11] by detecting the fracture region through the methods such as detector head decoupling, anchor-free, and augmentation strategies. They also showed that YOLOX can locate fractures even in low-featured X-ray images. Moon et al. [37] proposed a computer-aided facial bone fracture diagnosis (CA-FBFD) system based on the YOLOX model, effectively reducing the workload of doctors in diagnosing facial fractures in facial CT scans. While the application of the models of YOLO series to medical image recognition is a hot research topic, to date, no one has utilized YOLOv9 [26] for fracture detection.
@@
YOLOv8[16] 43.61 164.9 59 62.44 40.32 3.6
YOLOv8+SA[22] 43.64 165.4 62 63.99 41.49 3.9
YOLOv8+ECA[23] 43.64 165.5 61 62.64 40.21 3.6
YOLOv8+GAM[24] 49.29 183.5 60 63.32 40.74 8.7
YOLOv8+ResGAM[38] 49.29 183.5 62 63.97 41.18 9.4
YOLOv8+ResCBAM[25] 53.87 196.2 62 62.95 40.10 4.1
YOLOv9-C (Ours) 51.02 239.0 64 65.31 42.66 5.2
YOLOv9-E (Ours) 69.42 244.9 64 65.46 43.32 6.4
@@
3 Method
3.1 YOLOv9
3.1.1 Programmable Gradient Information
Programmable Gradient Information (PGI) is an auxiliary supervision framework designed to manage the propagation of gradient information across various semantic levels, to improve the detection capability of the model. PGI comprises three main components: main branch, auxiliary reversible branch, and multi-level auxiliary information. During the inference process, it exclusively employs the main branch, which handles both forward and back propagation. As the network becomes deeper, an information bottleneck may occur, leading to loss functions that fail to produce useful gradients. In such cases, auxiliary reversible branch employs reversible functions to preserve information integrity and mitigate information loss in the main branch. Additionally, multi-level auxiliary information addresses the issue of error accumulation from the deep supervision mechanism, improving the learning capacity of the model through the introduction of supplementary information at different levels. Notably, research [26] highlighted the efficacy of PGI in preserving information during training, particularly in scenarios with limited features. This provides the theoretical basis for the YOLOv9 model to have excellent performance in fracture detection tasks.
@@ 3.1.2 Generalized Efficient Layer Aggregation Network
3.2 Data Processing and Augmentation
Fig 2 illustrates the flowchart of the experiments conducted in this study. Since the publisher of the GRAZPEDWRI-DX [20] dataset did not provide predefined training, validation, and test sets, we randomly assigned 70% to training set, 20% to validation set, and 10% to test set during the data processing. Moreover, due to the limited brightness diversity of low-featured X-ray images, models trained only on these images may not generalize well to X-ray images in other environments. To enhance the robustness of the model, we employed data augmentation techniques to extend the training set. Specifically, we fine-tuned the contrast and luminance of the X-ray images using the addWeighted function from the OpenCV library.
@@ 3.2 Data Processing and Augmentation
YOLOv8[16] 43.61 164.9 62 63.63 40.41 7.7
YOLOv8+SA[22] 43.64 165.4 63 64.25 41.64 8.0
YOLOv8+ECA[23] 43.64 165.5 65 64.26 41.94 7.7
YOLOv8+GAM[24] 49.29 183.5 65 64.26 41.00 12.7
YOLOv8+ResGAM[38] 49.29 183.5 64 64.98 41.75 18.1
YOLOv8+ResCBAM[25] 53.87 196.2 64 65.78 42.16 8.7
YOLOv9-C (Ours) 51.02 239.0 66 65.57 43.70 12.7
YOLOv9-E (Ours) 69.42 244.9 66 65.62 43.73 16.1
@@ 4.1 Dataset
GRAZPEDWRI-DX [20] is a public dataset provided by the Medical University of Graz, which contains 20,327 X-ray images of pediatric wrist trauma. These X-ray images were collected by a team of pediatric radiologists at the University Hospital Graz from 2008 to 2018. The dataset comprises 6,091 patients and 10,643 studies, with a total of 74,459 labeled images, representing 67,771 labeled objects.
@@ 4.2 Experiment setup
The experiments in this paper utilized one single NVIDIA GeForce RTX 3090 GPU, employing Python with the PyTorch framework. Before training our model, we employed the YOLOv9 model weights pretrained on the MS COCO 2017 [27] dataset. In the training process, we trained the model using the SGD [41] optimizer, with a weight decay rate set to 5e-4 and a momentum of 0.937. We followed the research [21] to set the initial learning rate to 1e-2, the number of epochs to 100. Due to resource limitations (24GB memory) imposed by a single GPU, a batch size of 16 was employed for training the model.
@@ 4.3 Experimental Results
To evaluate the performance of YOLOv9 and other SOTA models in real diagnostic scenarios, this study compares model size (parameters and floating-point operations per second), accuracy (F1 score, mean average precision at 50% (mAP 50), and mean average precision from 50% to 95% (mAP 50-95)), and inference time. It is widely recognized that using larger input image sizes improves prediction accuracy but also requires more computational resources. Therefore, we conducted two experiments with input image sizes of 640 and 1024 for various scenarios, and the results are presented in Tables 1 and 2. With the input size of 640, both YOLOv9-C (Compact) and YOLOv9-E (Extended) demonstrate significantly improved mAP, while maintaining a reasonable inference speed. Specifically, YOLOv9-E achieves mAP 50-95 of 43.32%, which is 4.4% higher than 41.49% achieved by the current SOTA model YOLOv8+SA. When the input image size is 1024, the mAP 50-95 of YOLOv9-E reaches 43.73%, which also obtains the SOTA performance. However, due to the increased inference time, it is more suitable for deployment on devices with high computing resources.

### v2

@@ 1 Introduction
The accurate interpretation of X-ray images is critical to the success of surgery. However, there is a shortage of experienced radiologists in some areas, which poses a challenge [1, 2]. Computer-assisted diagnosis (CAD) helps specialists such as radiologists and surgeons to interpret medical images, including magnetic resonance imaging (MRI), computed tomography (CT), and X-ray images. However, it cannot be relied on as the sole means of detection due to uncertainties in its prediction accuracy. With the rapid development of deep learning in computer vision tasks [3, 4], its application to medical image processing has obtained satisfactory results [5].
You Only Look Once (YOLO) series [6, 7, 8, 9] are the main neural networks for real-time object detection task, widely employed in fracture detection [10, 11, 12]. Wrist fractures in children are more common cases and the GRAZPEDWRI-DX dataset [13] provides 20,327 X-ray images of pediatric wrist trauma that can be used in fracture detection tasks. Ju et al. [11] first used the YOLOv8 [8] model for fracture detection on this dataset. Since attention mechanisms [14, 15, 16, 17, 18] have excellent results in enhancing the performance of neural network models, Chien et al. [12] achieved the state-of-the-art (SOTA) performance by incorporating different attention mechanisms into the YOLOv8 model.
With the presentation of YOLOv9 [9], which achieved remarkable model performance on the MS COCO 2017 [19] benchmark dataset, achieving an impressive increase in mAP from 53.9% in YOLOv8 to 55.6%. This paper first trained the YOLOv9 model on the GRAZPEDWRI-DX dataset and obtained the SOTA performance, as shown in Fig 1.
@@
Figure 1: Comparisons of fracture detection models on the GRAZPEDWRI-DX dataset when the input image size is 640. In terms of accuracy, our models outperform all previous models at the state-of-the-art level.
@@ 2 Related Works
In the field of object detection task, detectors typically employ either one-stage or two-stage algorithms. Compared to two-stage object detectors, YOLO series models offer a balanced combination of accuracy and inference speed, making them suitable for deployment on mobile computing platforms for medical image recognition. Hržić et al. [10] employed the YOLOv4 [7] model for fracture detection on the GRAZPEDWRI-DX dataset [13], which was the first to demonstrate that the models of YOLO series can assist radiologists in more accurately predicting wrist injuries in children on X-ray images. Ju et al. [11] developed the “Fracture Detection Using YOLOv8 App” to aid surgeons in interpreting X-ray images for fractures, aiming to reduce misclassification and enhance the information available for fracture surgery. Chien et al. [12] proposed the YOLOv8-AM model, incorporating four attention modules [15, 16, 17, 18], and trained them on the GRAZPEDWRI-DX dataset [13]. Specifically, the ResCBAM-based YOLOv8-AM model achieved a significant improvement in mAP 50 value, increasing from 63.6% to 65.8% compared to the YOLOv8 model, which reached the performance of SOTA. While the application of the models of YOLO series to medical image recognition is a hot research topic, to date, no one has utilized YOLOv9 [9] for fracture detection.
@@
Figure 3: Visualization of the accuracy of predicting each class using different models on the GRAZPEDWRI-DX dataset with the input image size of 1024. (a) YOLOv8, (b) YOLOv8+ResCBAM (SOTA), (c) YOLOv9 (Ours).
3 Method
3.1 YOLOv9
3.1.1 Programmable Gradient Information
PGI is an auxiliary supervision framework designed to manage the propagation of gradient information across various semantic levels, to improve the detection capability of the model. PGI comprises three main components: main branch, auxiliary reversible branch, and multi-level auxiliary information. During the inference process, it exclusively employs the main branch, which handles both forward and back propagation. As the network becomes deeper, an information bottleneck may occur, leading to loss functions that fail to produce useful gradients. In such cases, auxiliary reversible branch employs reversible functions to preserve information integrity and mitigate information loss in the main branch. Additionally, multi-level auxiliary information addresses the issue of error accumulation from the deep supervision mechanism, improving the learning capacity of the model through the introduction of supplementary information at different levels. Notably, research [9] highlighted the efficacy of PGI in preserving information during training, particularly in scenarios with limited features. This provides the theoretical basis for the YOLOv9 model to have excellent performance in fracture detection tasks.
@@ 3.1.1 Programmable Gradient Information
YOLOv8[8] 43.61 164.9 0.59 62.44 40.32 3.6
YOLOv8+SA[17] 43.64 165.4 0.62 63.99 41.49 3.9
YOLOv8+ECA[16] 43.64 165.5 0.61 62.64 40.21 3.6
YOLOv8+GAM[18] 49.29 183.5 0.60 63.32 40.74 8.7
YOLOv8+ResGAM[12] 49.29 183.5 0.62 63.97 41.18 9.4
YOLOv8+ResCBAM[15] 53.87 196.2 0.62 62.95 40.10 4.1
YOLOv9-C (Ours) 51.02 239.0 0.64 65.31 42.66 5.2
YOLOv9-E (Ours) 69.42 244.9 0.64 65.46 43.32 6.4
@@
[…]
@@ 3.1.2 Generalized Efficient Layer Aggregation Network
YOLOv8[8] 43.61 164.9 0.62 63.63 40.41 7.7
YOLOv8+SA[17] 43.64 165.4 0.63 64.25 41.64 8.0
YOLOv8+ECA[16] 43.64 165.5 0.65 64.26 41.94 7.7
YOLOv8+GAM[18] 49.29 183.5 0.65 64.26 41.00 12.7
YOLOv8+ResGAM[12] 49.29 183.5 0.64 64.98 41.75 18.1
YOLOv8+ResCBAM[15] 53.87 196.2 0.64 65.78 42.16 8.7
YOLOv9-C (Ours) 51.02 239.0 0.66 65.57 43.70 12.7
YOLOv9-E (Ours) 69.42 244.9 0.66 65.62 43.73 16.1
@@
[…]
3.2 Data Processing and Augmentation
Fig 2 illustrates the flowchart of the experiments conducted in this study. Since the publisher of the GRAZPEDWRI-DX [13] dataset did not provide predefined training, validation, and test sets, we randomly assigned 70% to training set, 20% to validation set, and 10% to test set during the data processing. Moreover, due to the limited brightness diversity of low-featured X-ray images, models trained only on these images may not generalize well to X-ray images in other environments. To enhance the robustness [22, 23, 24] of the model, we followed the data augmentation technique used by Ju et al. [11] to extend the training set. Specifically, we fine-tuned the contrast and luminance of the X-ray images using the addWeighted function from the OpenCV library.
@@ 4.1 Dataset
GRAZPEDWRI-DX [13] is a public dataset provided by the Medical University of Graz, which contains 20,327 X-ray images of pediatric wrist trauma. These X-ray images were collected by a team of pediatric radiologists at the University Hospital Graz from 2008 to 2018. The dataset comprises 6,091 patients and 10,643 studies, with a total of 74,459 labeled images, representing 67,771 labeled objects. There is a significant class imbalance in this dataset. Specifically, the top two labeled objects are “text” (23,722 instances) and “fracture” (18,090 instances), whereas the number of instances for “bone anomaly” and “soft tissue” are only 276 and 464, respectively. All models use the same training data.
@@ 4.2 Experiment setup
The experiments in this paper utilized one single NVIDIA GeForce RTX 3090 GPU, employing Python with the PyTorch framework. Before training our model, we employed the YOLOv9 model weights pretrained on the MS COCO 2017 [19] dataset. In the training process, we trained the model using the SGD optimizer, with a weight decay rate set to 5e-4 and a momentum of 0.937. We followed the research [11] to set the initial learning rate to 1e-2, the number of epochs to 100. Due to resource limitations (24GB memory) imposed by a single GPU, a batch size of 16 was employed for training the model.
@@ 4.3 Experimental Results
It is well known that the YOLO series models are designed for real-time object detection and are suitable to be deployed in web or mobile applications to assist surgeons in pediatric wrist fracture diagnosis in all regions of the world. The research [13] has shown that the GRAZPEDWRI-DX dataset is the first publicly available and pertinent pediatric dataset, so we evaluate the performance of different models on it.
To evaluate the performance of YOLOv9 and other SOTA models in real diagnostic scenarios, this study compares model size (parameters and floating point operations), accuracy (F1 score, mean average precision at 50% (mAP 50), and mean average precision from 50% to 95% (mAP 50-95)), and inference time. It is widely recognized that using larger input image sizes improves prediction accuracy but also requires more computational resources. Therefore, we conducted two experiments with input image sizes of 640 and 1024 for various scenarios, and the results are presented in Tables 1 and 2. With the input size of 640, both YOLOv9-C (Compact) and YOLOv9-E (Extended) demonstrate significantly improved mAP, while maintaining a reasonable inference speed. Specifically, YOLOv9-E achieves mAP 50-95 of 43.32%, which is 4.4% higher than 41.49% achieved by the current SOTA model YOLOv8+SA. When the input image size is 1024, the mAP 50-95 of YOLOv9-E reaches 43.73%, which also obtains the SOTA performance. However, due to the increased inference time, it is more suitable for deployment on devices with high computing resources. In summary, the YOLOv9 model effectively balances inference time with fracture detection accuracy, making it suitable for deployment in web applications as a CAD system to assist surgeons in analyzing X-ray images.
To facilitate further performance enhancement, we visualized the accuracy of various models in predicting each class, as illustrated in Fig. 3. Across all models, the ability to accurately predict “fracture”, “metal”, and “text” classes exceeded 90%, with YOLOv9 notably achieving a 94.8% accuracy in “metal” class prediction. However, the performance in predicting bone anomalies and soft tissue classes was poorer, with YOLOv9 achieving only 11.3% and 28.2% accuracy, respectively. This significantly impacted the average accuracy of YOLOv9 in fracture detection.
5 Discussion
The performance of YOLO series models is significantly influenced by the quality and diversity of the training datasets. As mentioned above, the limited ability of these methods to accurately predict “bone anomaly” and “soft tissue” classes is due to the insufficient number of relevant data in the GRAZPEDWRI-DX dataset. Therefore, a future research can enhance performance by improving the models with additional information on “bone anomaly” and “soft tissue” classes.
This study focuses exclusively on the detection of pediatric wrist fractures, limiting its applicability to fractures in other body regions. To enhance the model’s utility across new cases, we have made the trained weight files available on our GitHub. Users can leverage these files to train the model using X-ray images of different fracture types, enabling the model to achieve reliable predictions with minimal training data.